OpenStudy (anonymous):

as a judge, judith makes a lot of money. She deposits 45,000.00 into an account that compounds its interest monthly. If her rate is 3.6%. How much interest will she receive after 10 years. If she doesnt make any withdraws or deposits. P= r= n= t=

OpenStudy (campbell_st):

is the interest rate per month or per year...?

OpenStudy (anonymous):

monthly

OpenStudy (campbell_st):

ok... then use \[A = P(1 + r)^t\] A = future value, P = principal, 45000 r = 0.036 t = 120 so \[A = 45000(1 + 0.036)^{120}\] to get the future value

OpenStudy (anonymous):

ok i appreciate all your help!!!